Given: angle DEG=(x+3y)degrees
angle GEF+(2x+y)degrees
angle DEF is a right angle
a. Solve for y in terms of x (what does this mean?) I have tried this what am I doing wrong?

(x+y) + (2x+y) = 90degrees
3x+4y+90
4/4y +90/4-3x
y=22.5-3

Answer by mananth(16946)

(x+y) + (2x+y) = 90
3x+2y=90
you have to make y the subject
add -3x to both sides
3x-3x+2y=90-3x
2y=90-3x
divide by 2
y=90-3x /2
y=45-(3/2)x